The sum of two consecutive odd integers equals their difference cubed. Find the smaller of the two.
Answer by drk(1908)

Let X = first number and X+2 = next odd integer.
(X + X + 2) = (x + 2 - x)^3
(2x+2) = 2^3
2x+2 = 8
2x = 6
X = 3.
the two numbers are 3 and 5.